OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -9 - 3 + 3 + 9 + ... + 81

OpenStudy (anonymous):

@amistre64 @ct3324432089

OpenStudy (anonymous):

ok wait let me check it

OpenStudy (anonymous):

first find out the no.of terms

OpenStudy (anonymous):

a1=-9,d=-3-(-9) = -3+9 =6(common difference) next tn = -9+(n-1)6 =-9-6+6n -15+6n now,6n-15 =81 6n=96 n=16 now s16= 16/2*(-9+81) =8* 72 = 576 look babes ,its done